OpenStudy (anonymous):

Which equation shows the substitution method being used to solve the system of linear equations? x + y = 6 x = y + 5 A. x = (x – 6) + 5 B. (y + 5) + y = 6 C. x + y = y + 5 D. x + (y + 5) = 6

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

Wel since we know that x = y + 5 we sub that into the first equation for 'x' \[\large (y + 5) + y = 6\]

OpenStudy (anonymous):

i thought it was B. well i acually knew it but i was making sure of it

OpenStudy (johnweldon1993):

Of course you did smarty pants ;P

OpenStudy (anonymous):

The system of equations is inconsistent. What are the missing values? 5x y=2 x+3y=8

OpenStudy (anonymous):

and im not a smarty apnts i hate that word

OpenStudy (anonymous):

pants*

OpenStudy (johnweldon1993):

Was there supposed to be a + or - sign in the first equation?

OpenStudy (anonymous):

+

OpenStudy (anonymous):

heeelloo....

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

@johnweldon1993